# Each of those eight coalition logics is also determined by six other kinds of models

### v1

@@ Abstract
Coalition Logic is a central logic in logical research on strategic reasoning. In two recent papers, Li and Ju argued that generally, concurrent game models, models of Coalition Logic, have three too strong assumptions: seriality, independence of agents, and determinism. They presented eight coalition logics based on eight classes of general concurrent game models, determined by which of the three assumptions they meet. In this paper, we show that each of the eight coalition logics is also determined by the following six kinds of models, with the respective properties: single-coalition-first action models; single-coalition-first neighborhood models; clear grand-coalition-first action models; clear single-coalition-first neighborhood models; tree-like grand-coalition-first action models; tree-like single-coalition-first neighborhood models.
@@ 1.1 Eight coalition logics determined by eight classes of general concurrent game models
Coalition Logic $\mathsf{CL}$ proposed by Pauly ([10, 11]) is a central logic in logical research on strategic reasoning.
@@
The first kind is concurrent game models. Roughly, in a concurrent game model: there are some states; there are some agents who form coalitions; for every coalition, there is an availability function, which specifies available joint actions of the coalition at states; for every coalition, there is an outcome function, which specifies possible outcome states of joint actions of the coalition.
The formula $\langle[\mathrm{C}]\rangle\phi$ is true at a state in a concurrent game model if $\mathrm{C}$ has an available joint action such that $\phi$ is true at every possible outcome state of the action in the model.
The second kind is alpha neighborhood models. Roughly, in an alpha neighborhood model: there are some states; there are some agents who form coalitions; for every coalition, there is an alpha neighborhood function, which specifies the alpha powers of the coalition. An alpha power is a scope into which the coalition can force the next moment to fall.
The formula $\langle[\mathrm{C}]\rangle\phi$ is true at a state in an alpha neighborhood model if $\mathrm{C}$ has an alpha power such that $\phi$ is true at every state of the power.
[9] proposed that concurrent game models have three too-strong assumptions: the independence of agents, seriality, and determinism. Based on general concurrent game models, which do not have the three assumptions, a Minimal Coalition Logic $\mathsf{MCL}$ was presented.
Although the three assumptions are generally too strong, we might want to keep some of them when constructing logics for strategic reasoning in some special kinds of situations. Considering which of the three properties we want to keep, there are eight coalition logics in total. [8] showed the completeness of the eight logics in a uniform way. It should be pointed out that the completeness $\mathsf{CL}$ is already given in [11, 5] and the completeness of $\mathsf{MCL}$ is already given in [9].
General concurrent game models are called grand-coalition-first action models in this paper.
@@ 1.2 Our work
In this work, we show that each of the eight coalition logics is also determined by six other kinds of models with respective properties.
  • The first kind is single-coalition-first action models. In this kind of models, the outcome function for single coalitions determines the outcome functions for other coalitions and the availability functions for all coalitions.
    This kind of models makes good sense for situations where we know the space of future possibilities and how agents’ behavior changes the future beforehand.
  • The second kind is single-coalition-first neighborhood models. In this kind of models, the neighborhood function for a coalition specifies the actual powers of the coalition. An actual power is a minimal scope in which the coalition can force the next moment to fall. In this kind of models, the neighborhood functions for single coalitions determine the neighborhood functions for other coalitions.
  • The third kind is clear grand-coalition-first action models. In this kind of models, different joint actions of the grand coalition do not have the same outcome.
  • The fourth kind is clear single-coalition-first neighborhood models. In this kind of models, the powers of an agent are pairwise disjoint.
  • The fifth kind is tree-like grand-coalition-first action models. In this kind of models, every state has a determined history.
  • The sixth kind is tree-like single-coalition-first neighborhood models. In this kind of models, every state has a determined history, too.
The rest of the paper is structured as follows.
  • In Section 2, we present some general settings of coalition logics, including their language, their action semantics and neighborhood semantics, and the transformation from the former to the latter.
  • In Section 3, we introduce the eight coalition logics determined by eight classes of grand-coalition-first action models.
  • In Section 4, we discuss compositionality of availability functions and outcome functions in action models.
  • In Section 5: we define single-coalition-first action models and show that they are grand-coalition-first action models; we define single-coalition-first neighborhood models, and show they can represent single-coalition-first action models.
  • In Section 6: we define clear grand-coalition-first action models, and show they are single-coalition-first action models; we define clear single-coalition-first neighborhood models, and show that they can represent clear grand-coalition-first action models.
  • In Section 7, we define tree-like grand-coalition-first action models, and show they are clear grand-coalition-first action models; we define tree-like single-coalition-first neighborhood models, and show they can represent tree-like grand-coalition-first action models.
  • In Section 8, we show that each of the eight coalition logics is determined by the six kinds of models with respective properties.
  • We make some concluding remarks in Section 9.
@@ 2.1 Language
Let $\mathtt{AP}$ be a countable set of atomic propositions and $\mathtt{AG}$ be a finite nonempty set of agents. The subsets of $\mathtt{AG}$ are called coalitions. $\mathtt{AG}$ is called the grand coalition.
@@
The formula $\langle[\mathrm{C}]\rangle\phi$ indicates that some available joint action of $\mathrm{C}$ ensures $\phi$. The propositional connectives ,$\bot$,$\lor$,$\rightarrow$ , ∨ , → and $\leftrightarrow$ are defined as usual. Define the dual $[\langle\mathrm{C}\rangle]\phi$ of $\langle[\mathrm{C}]\rangle\phi$ as $\neg\langle[\mathrm{C}]\rangle\neg\phi$, indicating every available joint action of $\mathrm{C}$ enables $\phi$.
@@ 2.2 Action semantics
Let $\mathtt{AC}$ be a nonempty set of actions. For any $\mathrm{C}\subseteq\mathtt{AG}$, we define $\mathtt{JA}_{\mathrm{C}}=\{\sigma_{\mathrm{C}}\mid\sigma_{\mathrm{C}}:\mathrm{C}\rightarrow\mathtt{AC}\}$, which is a set of joint actions of $\mathrm{C}$. For any $a\in\mathtt{AG}$, we write $\mathtt{JA}_{a}$ instead of $\mathtt{JA}_{\{a\}}$. We define $\mathtt{JA}=\bigcup\{\mathtt{JA}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$.
@@
  • for any $\mathrm{C}\subseteq\mathtt{AG}$, $\mathtt{av}_{\mathrm{C}}:\mathtt{ST}\rightarrow\mathcal{P}(\mathtt{JA}_{\mathrm{C}})$ is an available action function for $\mathrm{C}$.
  • for any $\mathrm{C}\subseteq\mathtt{AG}$, $\mathsf{out}_{\mathrm{C}}:\mathtt{ST}\times\mathtt{JA}_{\mathrm{C}}\rightarrow\mathcal{P}(\mathtt{ST})$ is an outcome function for $\mathrm{C}$.
@@
Different logics might put different constraints on action models.
Definition 3 (Action semantics of $\Phi$).
$\mathtt{AM},s\Vdash\top$ always
@@
It can be verified:
$\mathtt{AM},s\Vdash[\langle\mathrm{C}\rangle]\phi$Leftrightarrow for all $\sigma_{\mathrm{C}}\in\mathtt{av}_{\mathrm{C}}(s,\mathrm{C})$, there is $t\in\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})$ such that $\mathtt{AM},t\Vdash\phi$
@@ 2.3 Neighborhood semantics
A neighborhood model $\mathtt{NM}=(\mathtt{ST},\{\mathtt{nei}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ is called an alpha neighborhood model if for every $\mathrm{C}\subseteq\mathtt{AG}$ and $s\in\mathtt{ST}$, $\mathtt{nei}_{\mathrm{C}}(s)$ is closed under supersets.
Different logics may put different constraints on neighborhood models.
Definition 5 (Neighborhood semantics of $\Phi$).
$\mathtt{NM},s\Vdash\top$ always
@@
The following fact indicates that we can transform a pointed neighborhood model to a pointed alpha neighborhood model without changing which formulas it satisfies.
@@
Let $\mathtt{NM}=(\mathtt{ST},\{\mathtt{nei}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ be a neighborhood model and $\mathtt{NM}^{\alpha}=(\mathtt{ST},\mathtt{AG},$$\{\mathtt{nei}^{\alpha}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ be an alpha neighborhood model such that for every $\mathrm{C}\subseteq\mathtt{AG}$ and $s\in\mathtt{ST}$, $\mathtt{nei}^{\alpha}_{\mathrm{C}}(s)$ is the closure of $\mathtt{nei}_{\mathrm{C}}(s)$ under supersets.
@@ 2.4 Transformation of action semantics to neighborhood semantics
Definition 6 (Actual and alpha effectivity functions of action models).
Let $\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ be an action model, and $\mathrm{C}\subseteq\mathtt{AG}$.
The function $\mathtt{AE}_{\mathrm{C}}$ defined as follows is called the actual effectivity function for $\mathrm{C}$ of $\mathtt{AM}$: for every $s\in\mathtt{ST}$,
@@
The function $\mathtt{LE}_{\mathrm{C}}$ defined as follows is called the alpha effectivity function for $\mathrm{C}$ of $\mathtt{AM}$: for every $s\in\mathtt{ST}$,
@@
In a sense, actual effectivity functions and alpha effectivity functions of action models contain important information on coalition powers.
@@
We say a neighborhood model $\mathtt{NM}=(\mathtt{ST},\{\mathtt{nei}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ $z$-represents an action model $\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ if for every $\mathrm{C}\subseteq\mathtt{AG}$, $\mathtt{nei}_{\mathrm{C}}=\mathtt{AE}_{\mathrm{C}}$.
We say a class of neighborhood models $\mathbf{NM}$ $z$-represents a class of action models $\mathbf{AM}$ if (1) every action model in $\mathbf{AM}$ is $z$-representable by a neighborhood mode in $\mathbf{NM}$, and (2) every neighborhood model in $\mathbf{NM}$ $z$-represents an action mode in $\mathbf{AM}$.
We say an alpha neighborhood model $\mathtt{NM}=(\mathtt{ST},\{\mathtt{nei}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ $\alpha$-represents an action model $\mathtt{AM}$ if for every $\mathrm{C}\subseteq\mathtt{AG}$, $\mathtt{nei}_{\mathrm{C}}=\mathtt{LE}_{\mathrm{C}}$.
We say a class of alpha neighborhood models $\mathbf{NM}$ $\alpha$-represents a class of action models $\mathbf{AM}$ if (1) every action model in $\mathbf{AM}$ is $\alpha$-representable by an alpha neighborhood model in $\mathbf{NM}$, and (2) every alpha neighborhood model in $\mathbf{NM}$ $\alpha$-represents an action mode in $\mathbf{AM}$.
@@
    For every action model $\mathtt{AM}$ and neighborhood model $\mathtt{NM}$, if $\mathtt{AM}$ is $z$-representable by $\mathtt{NM}$, then for every $s$ of $\mathtt{AM}$ and $\phi$ in $\Phi$, $\mathtt{AM},s\Vdash\phi$ if and only if $\mathtt{NM},s\Vdash\phi$.
    Consequently, for every class of action models $\mathbf{AM}$ and class of neighborhood models $\mathbf{NM}$, if $\mathbf{AM}$ is $z$-representable by $\mathbf{NM}$, then for every $\phi$ in $\Phi$, $\phi$ is valid with respect to $\mathbf{AM}$ if and only if $\phi$ is valid with respect to $\mathbf{NM}$.
@@
    For every action model $\mathtt{AM}$ and alpha neighborhood model $\mathtt{NM}$, if $\mathtt{AM}$ is $\alpha$-representable by $\mathtt{NM}$, then for every $s$ of $\mathtt{AM}$ and $\phi$ in $\Phi$, $\mathtt{AM},s\Vdash\phi$ if and only if $\mathtt{NM},s\Vdash\phi$.
    Consequently, for every class of action models $\mathbf{AM}$ and class of alpha neighborhood model $\mathbf{NM}$, if $\mathbf{AM}$ is $\alpha$-representable by $\mathbf{NM}$, then for every $\phi$ in $\Phi$, $\phi$ is valid with respect to $\mathbf{AM}$ if and only if $\phi$ is valid with respect to $\mathbf{NM}$.
Proof.
Here, we show only (1). We can show (2) in a similar way.
Let $\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ be an action model and $\mathtt{NM}=(\mathtt{ST},\{\mathtt{nei}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ be a neighborhood model such that $\mathtt{AM}$ is $z$-representable by $\mathtt{NM}$. Let $s\in\mathtt{ST}$ and $\phi\in\Phi$. We want to show $\mathtt{AM},s\Vdash\phi$ if and only if $\mathtt{NM},s\Vdash\phi$.
The proof is by induction of $\phi$.
It is trivial when $\phi$ is a proposition letter p, $\phi$ is $\top$, $\phi$ is $\neg\psi$ and $\phi$ is $\psi\wedge\theta$. We are done if we can show when $\phi$ is $\langle[\mathrm{C}]\rangle\psi$, $\mathtt{AM},s\Vdash\phi$ if and only if $\mathtt{NM},s\Vdash\phi$.
$\Rightarrow$
Assume $\mathtt{AM},s\Vdash\phi$. Then, $\mathtt{AM},s\Vdash\langle[\mathrm{C}]\rangle\psi$. Then, there is $\sigma_{{\mathrm{C}}}\in\mathtt{av}_{\mathrm{C}}(s)$ such that for all $t\in\mathsf{out}_{\mathrm{C}}(s,\sigma_{{\mathrm{C}}})$, $\mathtt{AM},t\Vdash\psi$. Note $\mathtt{nei}_{\mathrm{C}}=\mathtt{AE}_{\mathrm{C}}$. Then, there is $\mathrm{Y}\in\mathtt{nei}_{\mathrm{C}}(s)$ such that $\mathrm{Y}=\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})$. Then, for all $t\in\mathrm{Y}$, $\mathtt{NM},t\Vdash\psi$. Then, $\mathtt{NM},s\Vdash\phi$.
$\Leftarrow$
Assume $\mathtt{NM},s\Vdash\phi$. Then, $\mathtt{NM},s\Vdash\langle[\mathrm{C}]\rangle\psi$. Then, there is $\mathrm{Y}\in\mathtt{nei}_{\mathrm{C}}(s)$ such that for all $t\in\mathrm{Y}$, $\mathtt{NM},t\Vdash\psi$. Note $\mathtt{nei}_{\mathrm{C}}=\mathtt{AE}_{\mathrm{C}}$. Then, $\mathrm{Y}\in\mathtt{AE}_{\mathrm{C}}(s)$. Then, there is $\sigma_{\mathrm{C}}\in\mathtt{av}_{\mathrm{C}}(s)$ such that for all $t\in\mathsf{out}_{\mathrm{C}}(s,\sigma_{{\mathrm{C}}})$, $\mathtt{AM},t\Vdash\psi$. Then, $\mathtt{AM},s\Vdash\phi$.

### v2

@@ Abstract
Coalition Logic is a central logic in logical research on strategic reasoning. In two recent papers, Li and Ju argued that generally, concurrent game models, models of Coalition Logic, have three too strong assumptions: seriality, independence of agents, and determinism. They presented eight coalition logics based on eight classes of general concurrent game models, determined by which of the three assumptions are met. In this paper, we show that each of the eight logics is also determined by six other kinds of models, with respective properties, that is, single-coalition-first action models, single-coalition-first neighborhood models, clear grand-coalition-first action models, clear single-coalition-first neighborhood models, tree-like grand-coalition-first action models, and tree-like single-coalition-first neighborhood models.
@@ 1.1 Eight coalition logics determined by eight classes of general concurrent game models
Coalition Logic $\mathsf{CL}$, proposed by Pauly ([10, 11]), is a central logic in logical research on strategic reasoning.
@@
The first kind is concurrent game models. Roughly, in a concurrent game model: there are some states; there are some agents, who can form coalitions; for every coalition, there is an availability function, specifying available joint actions of the coalition at states; for every coalition, there is an outcome function, specifying possible outcome states of joint actions of the coalition. The formula $\langle[\mathrm{C}]\rangle\phi$ is true at a state in a concurrent game model if $\mathrm{C}$ has an available joint action such that $\phi$ is true at every possible outcome state of the action.
The second kind is alpha neighborhood models. Alpha neighborhood models do not have actions. Consequently, they do not have availability functions and outcome functions. Instead, for every coalition, there is a neighborhood function that specifies the alpha powers of the coalition. An alpha power is a set of states into which the coalition can force the next moment to fall. The formula $\langle[\mathrm{C}]\rangle\phi$ is true at a state in an alpha neighborhood model if $\mathrm{C}$ has an alpha power such that $\phi$ is true at every state of the power.
[9] argued that concurrent game models have three too strong assumptions: seriality, the independence of agents, and determinism. Based on general concurrent game models, which do not have the three assumptions, they proposed a Minimal Coalition Logic $\mathsf{MCL}$. Although the three assumptions are generally too strong, we might want to keep some of them when constructing logics for strategic reasoning in some kinds of situations. Considering which of the three properties we want to keep, there are eight coalition logics in total. [8] showed the completeness of the eight logics in a uniform way.
@@ 1.2 Our work
In this paper, we call general concurrent game models grand-coalition-first action models. In this work, we show that each of the eight coalition logics is also determined by six other kinds of models, with respective properties.
The first kind is single-coalition-first action models. In this kind of models, the outcome functions for single coalitions determine the outcome functions for other coalitions. This kind of models makes good sense for situations where every agent can change a component of the state, and different components of the state that can be changed by different agents are independent of each other. The second kind is single-coalition-first neighborhood models. In this kind of models, the neighborhood function for a coalition specifies the actual powers of the coalition. An actual power is a minimal scope into which the coalition can force the next moment to fall. In this kind of models, the neighborhood functions for single coalitions determine the neighborhood functions for other coalitions.
The third kind is clear grand-coalition-first action models. In this kind of models, different joint actions of the grand coalition have different outcomes. Clear grand-coalition-first action models are quite common in game theory. The fourth kind is clear single-coalition-first neighborhood models. In this kind of models, the actual powers of an agent are pairwise disjoint.
The fifth kind is tree-like grand-coalition-first action models. In this kind of models, every state has a determined history. The sixth kind is tree-like single-coalition-first neighborhood models. In this kind of models, every state has a determined history, too.
The rest of the paper is structured as follows. In Section 2, we present some general settings of coalition logics, including their language, their action semantics and neighborhood semantics, and the transformation from action semantics to neighborhood semantics. In Section 3, we introduce the eight coalition logics determined by eight classes of grand-coalition-first action models. In Section 4, we define single-coalition-first action models and show that they are grand-coalition-first action models; we define single-coalition-first neighborhood models and show that they can represent single-coalition-first action models. In Section 5, we define clear grand-coalition-first action models and show they are single-coalition-first action models; we define clear single-coalition-first neighborhood models and show that they can represent clear grand-coalition-first action models. In Section 6, we define tree-like grand-coalition-first action models and show that they are clear grand-coalition-first action models; we define tree-like single-coalition-first neighborhood models and show that they can represent tree-like grand-coalition-first action models. In Section 7, we show that each of the eight coalition logics is determined by the six kinds of models with respective properties. We conclude in Section 8.
@@ 2.1 Language
Let $\mathtt{AP}$ be a countable set of atomic propositions, and $\mathtt{AG}$ be a finite nonempty set of agents. The subsets of $\mathtt{AG}$ are called coalitions, and $\mathtt{AG}$ is called the grand coalition. In the sequel, when no confusion arises, we often write $a$ instead of $\{a\}$, where $a\in\mathtt{AG}$.
@@
The formula $\langle[\mathrm{C}]\rangle\phi$ indicates some available joint action of $\mathrm{C}$ ensures $\phi$. The propositional connectives ,$\bot$,$\lor$,$\rightarrow$ , ∨ , → and $\leftrightarrow$ are defined as usual.
@@ 2.2 Action semantics
Let $\mathtt{AC}$ be a nonempty set of actions. For every $\mathrm{C}\subseteq\mathtt{AG}$, we define $\mathtt{JA}_{\mathrm{C}}=\{\sigma_{\mathrm{C}}\mid\sigma_{\mathrm{C}}:\mathrm{C}\rightarrow\mathtt{AC}\}$, which is a set of joint actions of $\mathrm{C}$. Joint actions of the grand coalition are called action profiles. In the sequel, we sometimes use sequences of actions to represent joint actions of coalitions, where an order among agents is implicitly presupposed. We define $\mathtt{JA}=\bigcup\{\mathtt{JA}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$.
For every $\mathrm{C},\mathrm{D}\subseteq\mathtt{AG}$ such that $\mathrm{D}\subseteq\mathrm{C}$ and $\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$, we use $\sigma_{\mathrm{C}}|_{\mathrm{D}}$ to indicate the restriction of $\sigma_{\mathrm{C}}$ to $\mathrm{D}$, which is in $\mathtt{JA}_{\mathrm{D}}$.
@@
  • for every $\mathrm{C}\subseteq\mathtt{AG}$, $\mathtt{av}_{\mathrm{C}}:\mathtt{ST}\rightarrow\mathcal{P}(\mathtt{JA}_{\mathrm{C}})$ is an availability function for $\mathrm{C}$.
  • for every $\mathrm{C}\subseteq\mathtt{AG}$, $\mathsf{out}_{\mathrm{C}}:\mathtt{ST}\times\mathtt{JA}_{\mathrm{C}}\rightarrow\mathcal{P}(\mathtt{ST})$ is an outcome function for $\mathrm{C}$.
@@
Different coalition logics might put different constraints on action models.
Definition 3 (Action semantics).
$\mathtt{AM},s\Vdash\top$
@@ 2.3 Neighborhood semantics
A neighborhood model $\mathtt{NM}^{\alpha}=(\mathtt{ST},\{\mathtt{nei}^{\alpha}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ is an alpha neighborhood model if for every $\mathrm{C}\subseteq\mathtt{AG}$ and $s\in\mathtt{ST}$, $\mathtt{nei}^{\alpha}_{\mathrm{C}}(s)$ is closed under supersets.
Different coalition logics might put different constraints on neighborhood models.
Definition 5 (Neighborhood semantics).
$\mathtt{NM},s\Vdash\top$
@@
The following fact, which is easy to verify, indicates that we can transform a pointed neighborhood model to a pointed alpha neighborhood model without changing the set of formulas it satisfies.
@@
Let $\mathtt{NM}=(\mathtt{ST},\{\mathtt{nei}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ be a neighborhood model and $\mathtt{NM}^{\alpha}=(\mathtt{ST},\{\mathtt{nei}^{\alpha}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ be an alpha neighborhood model such that for all $\mathrm{C}\subseteq\mathtt{AG}$ and $s\in\mathtt{ST}$, $\mathtt{nei}^{\alpha}_{\mathrm{C}}(s)$ is the closure of $\mathtt{nei}_{\mathrm{C}}(s)$ under supersets.
@@ 2.4 Transformation of action semantics to neighborhood semantics
Definition 6 (Actual effectivity functions and alpha effectivity functions of action models).
Let $\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ be an action model.
For every $\mathrm{C}\subseteq\mathtt{AG}$, the function $\mathtt{AE}_{\mathrm{C}}$ defined as follows is called the actual effectivity function for $\mathrm{C}$ in $\mathtt{AM}$: for every $s\in\mathtt{ST}$,
@@
For every $\mathrm{C}\subseteq\mathtt{AG}$, the function $\mathtt{LE}_{\mathrm{C}}$ defined as follows is called the alpha effectivity function for $\mathrm{C}$ in $\mathtt{AM}$: for every $s\in\mathtt{ST}$,
@@
An action model $\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ is $z$-representable by a neighborhood model $\mathtt{NM}=(\mathtt{ST},\{\mathtt{nei}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ if for every $\mathrm{C}\subseteq\mathtt{AG}$, $\mathtt{AE}_{\mathrm{C}}=\mathtt{nei}_{\mathrm{C}}$.
A class of action models $\mathbf{AM}$ is $z$-representable by a class of neighborhood models $\mathbf{NM}$ if (1) every action model in $\mathbf{AM}$ is $z$-representable by a neighborhood mode in $\mathbf{NM}$, and (2) every neighborhood model in $\mathbf{NM}$ $z$-represents an action mode in $\mathbf{AM}$.
An action model $\mathtt{AM}$ is $\alpha$-representable by an alpha neighborhood model $\mathtt{NM}^{\alpha}=(\mathtt{ST},\{\mathtt{nei}^{\alpha}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ if for every $\mathrm{C}\subseteq\mathtt{AG}$, $\mathtt{LE}_{\mathrm{C}}=\mathtt{nei}^{\alpha}_{\mathrm{C}}$.
A class of action models $\mathbf{AM}$ is $\alpha$-representable by a class of alpha neighborhood models $\mathbf{NM}^{\alpha}$ if (1) every action model in $\mathbf{AM}$ is $\alpha$-representable by an alpha neighborhood model in $\mathbf{NM}^{\alpha}$, and (2) every alpha neighborhood model in $\mathbf{NM}^{\alpha}$ $\alpha$-represents an action mode in $\mathbf{AM}$.
@@
    For every action model $\mathtt{AM}$ and neighborhood model $\mathtt{NM}$, if $\mathtt{AM}$ is $z$-representable by $\mathtt{NM}$, then for every state $s$ of $\mathtt{AM}$ and formula $\phi$ in $\Phi$, $\mathtt{AM},s\Vdash\phi$ if and only if $\mathtt{NM},s\Vdash\phi$.
    Consequently, for every class of action models $\mathbf{AM}$ and class of neighborhood models $\mathbf{NM}$, if $\mathbf{AM}$ is $z$-representable by $\mathbf{NM}$, then for every formula $\phi$ in $\Phi$, $\phi$ is valid with respect to $\mathbf{AM}$ if and only if $\phi$ is valid with respect to $\mathbf{NM}$.
@@
    For every action model $\mathtt{AM}$ and alpha neighborhood model $\mathtt{NM}^{\alpha}$, if $\mathtt{AM}$ is $\alpha$-representable by $\mathtt{NM}^{\alpha}$, then for every state $s$ of $\mathtt{AM}$ and formula $\phi$ in $\Phi$, $\mathtt{AM},s\Vdash\phi$ if and only if $\mathtt{NM}^{\alpha},s\Vdash\phi$.
    Consequently, for every class of action models $\mathbf{AM}$ and class of alpha neighborhood model $\mathbf{NM}^{\alpha}$, if $\mathbf{AM}$ is $\alpha$-representable by $\mathbf{NM}^{\alpha}$, then for every formula $\phi$ in $\Phi$, $\phi$ is valid with respect to $\mathbf{AM}$ if and only if $\phi$ is valid with respect to $\mathbf{NM}^{\alpha}$.